### v3

# QF-tuner: Breaking Tradition in Reinforcement Learning

Abstract
Hyperparameter tuning in reinforcement learning algorithms refers to choosing the optimal parameters that may increase the algorithm’s performance. Manual or random hyperparameter tuning methods can be problematic, as even slight variations in their values can result in significantly different outcomes in the learning process. In this paper, we propose a new method, QF-tuner, for automatic hyperparameter tuning in the Q-learning algorithm using the FOX optimization algorithm (FOX). A new objective function has been proposed for the FOX, prioritizing reward over learning error and time. QF-tuner starts by running the FOX and tries to minimize the fitness value derived from observations at each iteration by executing the Q-learning algorithm. The proposed method has been evaluated using two control tasks from the OpenAI Gym: CartPole and FrozenLake. The empirical results of the QF-tuner on the CartPole control task show a reward of 499, and on the FrozenLake control task, a reward of 1. These results indicate that the QF-tuner outperforms other optimization algorithms. On the FrozenLake control task, there was a 36% increase in reward with a 26% reduction in learning time; on the CartPole control task, there was a 57% increase in reward with a 20% decrease in learning time. Thus, the QF-tuner is an essential method for hyperparameter tuning in reinforcement learning algorithms, enabling more effective solutions to control task problems.
Keywords FOX Optimization Algorithm  $\cdot$ Hyperparameter  $\cdot$ Optimization  $\cdot$ QF-tuner  $\cdot$ Q-learning  $\cdot$ Reinforcement Learning.
1 Introduction
Reinforcement learning (RL) involves selecting the actions that achieve the maximum rewards in each state of the problem’s environment. This behavior is learned through interactions with the environment by observing its trial-and-error responses. It is similar to how children explore their surroundings and discover the actions that lead to goal achievement [1]. Q-learning is considered a standard algorithm in RL, which is model-free off-policy, widely used, and simple to implement to solve finite markov decision processess (MDPs). RL algorithms solve problems, such as gaming, controlling, and scheduling, by moving from one state to another based on an approximating action-value function in the problem’s environment to select the optimal action that make a correct decision in the current state [2, 3]. The process of approximating the action-value function is repeated several episodes to increase the reward and find the optimal policy [4]. The FOX optimization algorithm (FOX) is a recently proposed metaheuristic optimization algorithm that takes inspiration from nature by mimicking how foxes hunt prey [5]. However, in RL algorithms, hyperparameters are set before learning that controls the learning process. They affect how the algorithm explores and learns from the environment, such as the learning rate, discount factor, and exploration strategy [6, 7]. While RL algorithms can yield maximum rewards; they require overhead tuning the hyperparameter to achieve optimum performance [8]. The difficulty is in determining the most effective values for each hyperparameter, as even minor adjustments might influence the learning process in any RL algorithm. However, manual hyperparameter tuning can be complex and time-intensive. Several methods have been suggested for hyperparameter tuning manually, randomly, or semi-automatic adjusting values for each parameter [9]. Nevertheless, these methods may not generally apply to all RL problems due to environmental variables and conditions variations.
This paper employed the FOX for automatic hyperparameter tuning in the Q-learning algorithm to find the optimal parameter that results in maximum rewards and the lowest learning time (steps). This method, QF-tuner, solves the problems associated with manual hyperparameter tuning in RL algorithms. It offers a robust and effective solution that addresses the challenges of hyperparameter tuning. Furthermore, a new objective function has been proposed and utilized in the FOX. it includes reward, error, and learning time and prioritizes reward over error and learning time. However, the contributions of this paper are as follows:
  1. 1.
    QF-tuner uses the state-of-the-art FOX optimization algorithm for automatic hyperparameter tuning in Q-learning, addressing the complexities of manual or random tuning.
  2. 2.
    A new objective function is proposed that prioritizes reward over error and learning time.
  3. 3.
    Empirical results show that QF-tuner improves Q-learning performance, achieving higher rewards with fewer learning steps in control tasks like FrozenLake and CartPole.
QF-tuner is evaluated using two control tasks from OpenAI Gym: FrozenLake and CartPole. FrozenLake presents a grid-world design where the agent must navigate slippery terrain to reach a goal while avoiding hazards, requiring decision-making in a stochastic environment. A CartPole balances a pole on a moving cart with precise control to maintain stability in a changing environment [10]. These tasks are benchmarks for evaluating RL algorithms, offering many challenges that test the robustness of the proposed approach across different control situations. The empirical results show that QF-tuner has improved the overall Q-learning algorithm’s performance by increasing the rewards with minimum steps.
This paper is structured as follows: section 2 reviews the latest studies in hyperparameter tuning using optimization algorithms. section 3 discuss the OpenAI Gym control tasks and explain the Q-learning, FOX, and the QF-tuner in detail. section 4 list and discuss the experimental results. Finally, section 5 presents the conclusion and future work suggestions.
2 Related Work
In the past few years, many investigations have focused on solving control tasks using RL, including the conventional or improved Q-learning algorithm. The authors of [11] proposed a framework for reducing the burden of hyperparameter tuning by using random feature approximation to efficiently compute state features and a factorial policy that allows fast processing of many discrete actions. The proposed framework was able to reduce the overhead of hyperparameter tuning in the context of large-scale control tasks. Furthermore, the authors [12] suggested two Q-learning-based methods for solving the CartPole control task. They used mean squared error (MSE) and mean absolute error (MAE) as reward objective functions. The evaluation focused on how well each method balanced the pole, intending to maximize the reward. Their results show that Q-learning with MSE tends to ignore outliers while with MSE gives more weight to outlier values. Moreover, the authors [13] designed a software package with a scalable GA to explore the best hyperparameters in deep RL. This package has been tested on OpenAI Gym environments and compared to the Bayesian approach. The study’s conclusions show that this approach is effective at maximizing reward with fewer episodes
The authors [14] proposed a hybrid model combining the bees algorithm (BA) with Q-learning to improve feature selection in coronary heart disease diagnosis. The proposed method was experimented on multiple datasets and showed improved accuracy compared to conventional algorithms. Another study by [15] presented a comparison study on the FrozenLake control task between the Q-learning algorithm and cognitive instance-based learning (IBL) algorithms. The Q-learning and  IBL agents were tested on different grid sizes of Frozen Lake. The IBL algorithm surpassed the Q-learning algorithm by demanding less learning time and better generalizing. However, the study did not address Q-learning’s limitations, underscoring the challenge of developing an enhanced version capable of effectively solving FrozenLake and other control tasks. Moreover, the authors [16] suggested Convergence-AwarE SAmpling with scReening (CAESAR) aggregation scheme. CAESAR blends convergence-aware sampling with a screening mechanism to integrate knowledge from agents learning in similar MDPs, improving overall learning efficiency. The results of the Frozen Lake control task validate CAESAR’s significance in improving agent learning across various control task environments.
However, existing hyperparameter tuning methods like particle swarm optimization (PSO)genetic algorithms (GA), and BA often struggle to escape local optima in complex control tasks [14, 17, 18]. These methods rely on fixed or semi-adaptive strategies, limiting their effectiveness. In contrast, the QF-tuner introduces an automatic tuning method that adjusts hyperparameters based on novel objective functions, overcoming these limitations and improving stability and efficiency, especially in tasks where other algorithms face convergence issues.
3 Materials and Methods
This section presents materials (control tasks) and the proposed method (QF-tuner). It starts by explaining the OpenAI Gym environment control tasks—FrozenLake and CartPole. Subsequent sections will provide a detailed background of the Q-learning algorithm, FOX, and the QF-tuner.
3.1 Materials
This paper used two control tasks—FrozenLake and CartPole—developed by OpenAI Gym to evaluate the QF-tuner. In the FrozenLake control task, an agent has to cross a grid world over a FrozenLake. This task introduces obstacles, creating a challenging environment [19]. Furthermore, in the CartPole environment, an inverted pendulum on a moving cart needs to be stabilized. This task is common when evaluating the ability of RL algorithms to maintain stability in a continuous state space [20]. However, QF-tuner uses these tasks to guarantee a standardized evaluation framework for evaluating the performance. Moreover, this paper advances experiential and makes relevant comparisons with previous research findings using other methods on these control task environments.
3.2 Q-learning Algorithm
One of the most promising fields of artificial intelligence (AI) is RL, which can learn directly from interactions between the agent and the environment [21] .RL does not require learning data or a target label; it only needs to take action in the problem’s environment by the agent and attempt to earn a reward [22]. The agent’s learning process mimics how humans play a specific video game. The agent always tries to win; if it loses, it will avoid the behavior that led to the loss in the next attempt [23]. In 1989, Watkins introduced a method considered an early breakthrough in RL known as Q-learning, defined by the following equations [24]:
$\delta_{t}=R_{t+1}+\gamma\max_{a}Q(S_{t+1},a)-Q(S_{t},A_{t})$ (1)
$Q(S_{t},A_{t})=Q(S_{t},A_{t})+\alpha\delta_{t}$ (2)
Where $\delta$ is the temporal difference error, $R$ is the reward, $\alpha$ is the step size, $\gamma$ is the discount factor, $S_{t}$ is the current state, $S_{t+1}$ is the next state, $A_{t}$ is the current action, and $a$ is the selected action. The $\max_{a}$ ensures the use of the optimal Q-value of the next state. Moreover, $Q(S_{t},A_{t})$ is the action-value function representing the cumulative reward expected if action $A$ is taken in the state $S$. Action-value pairs are stored in a Q-table table with dimensions of [number of states, number of actions].
In the learning process, the Q-learning’s agent obtains the reward $R$ after taking the action $A$ in the current state $S$, then moves to the next state $S_{t+1}$. Additionally, the Q-learning algorithm tries to find the best course of actions that follow policy in the MDP by determining the optimal function of action-value using the update rule from eqs. 1 and 2 [25, 26, 27]. Moreover, the agent interacts with the environment to observe the reward and updates the action-value function to improve its policy estimates over time. Agent interactions process iteratively, allowing the algorithm to modify its policy based on observations of the problem’s environment. Finally, the agent should have the ability to select the best action (optimal policy) in a given state using the estimated Q-table by the following eq. 3 [28]:
$\pi^{*}=\arg\max_{a}Q^{*}(s,a)$ (3)
where $\pi^{*}$ represents the optimal policy and $\arg\max_{a}$ is a function returns an action that maximizes the value of $Q^{*}(a)$ based on learned policy [28, 29].
3.3 FOX optimization algorithm
FOX is a new optimization algorithm inspired by red foxes’ hunting behavior that aims to search for the the best solution with the best fitness value [5]. FOX has many search agents working together iteratively, each trying to find the best fitness value during the search. It involves two major phases: exploration and exploitation. In exploration, the Fox agent uses a random walk strategy to locate and catch prey, utilizing its ability to detect ultrasound. During its search for prey, the Fox agent may listen for sounds indicating where its prey is. In exploitation, the Fox agent can hear ultrasound, so it takes time for the sound of its prey to reach it. The speed of sound is 343 meters per second [30]. However, since this speed is always the same, FOX uses a different method to measure it. The Fox agent jumps when it thinks it can catch its prey based on how long the sound takes to reach it. Thus, the Fox agent’s ability to catch prey depends on the time it takes for the sound to reach them while jumping [31]. The following equations described the primary steps of FOX in detail.
$S=\frac{\text{Best}X}{T_{i}}$ (4)
$DSF_{i}=S\times T_{i}$ (5)
The variable $DSF_{i}$ represents the sound distance from the Fox agent. The sound speed in the medium is represented by the variable $S$ and calculated by eq. 4. The variable $T$ represents an arbitrary number in the $[0,1]$ range. The variable $i$ represents the current iteration of the optimization process [32]. Moreover, The distance of a Fox agent from its prey is defined as follows [33]:
$DFP_{i}=DSF_{i}\times 0.5$ (6)
Once a Fox agent has determined the distance between itself and its prey, it should jump and catch the prey accurately. The Fox agent must calculate the required jump height using the eq. 7 to perform this jump [5, 34].
$Jump_{i}=0.5\times 9.81\times t^{2}$ (7)
where 9.81 represents the acceleration due to gravity, and $t$ is the time average taken by sound to travel [5].
$X_{i+1}=DFP_{i}\times Jump_{i}\times c_{1}$ (8)
$X_{i+1}=DFP_{i}\times Jump_{i}\times c_{2}$ (9)
where the value of $c_{1}$ and $c_{2}$ are 0.180 and 0.820, respectively. These values have been determined based on the Fox agent’s jump. The direction of the jump is randomly chosen to be either northeast or opposite direction. However, the Fox agent continuously moves to a new position in exploration and exploitation. In exploitation, the new position is determined using either eq. 8 or eq. 9.
On the other hand, the exploration determined the new position using the following equation [5]:
$X_{i+1}=\text{Best}X\times\text{rnd}(1,\text{dim})\times\min(tt)\times a$ (10)
The variable $tt$ is the time average, which is equal to the summation of the variable $T$ used in eq. 4 divided by the dimension of the problem (dim). The exploration is controlled by variable $a$ that is computed by $2\times[i-(i/max(iter))]$. However, the source code of FOX implementation can be accessed by https://github.com/Hardi-Mohammed/FOX. The next subsequent section will present an explanation of the QF-tuner.
3.4 Hyperparameter Tuning Method (QF-tuner)
The Q-learning algorithm has several hyperparameters that affect the agent’s performance and behavior regarding the learned policy. From eqs. 2 and 1, the following hyperparameters can be noticed: step size ($\alpha$) and discount factor ($\gamma$). They should be set before the beginning of the learning process. The $\alpha$ is crucial in updating the action-values function at each time step. It determines the extent to which new information overrides old information, and its value is typically within the range $[0,1]$. A lower $\alpha$ value means that the algorithm will be less sensitive to new information and will take longer to converge, while a higher value results in faster learning but may lead to instability. Therefore, choosing the correct value of $\alpha$ is essential for achieving the best performance in any learning algorithm [35]. In comparison, the $\gamma$ affects future reward for an agent’s decisions and its value range $[0,1]$. A more significant value (close to 1) indicates that the agent emphasizes future rewards, while a smaller value means that the agent gives less importance to future rewards. This parameter controls the balance between immediate and future rewards, affecting the agent’s decision-making process [36].
Developing a multi-objective function was necessary to address the multiple objectives of increasing reward, reducing the learning error, and minimizing the learning time (steps). This function is needed to encapsulate the simultaneous optimization of these three objectives within a single fitness function. Thus, this paper suggested the following objective function:
$\textit{fitness}=\sum_{i=1}^{n}(2R-e)\times\frac{1}{st}$ (11)
where the $e$ variable represents the temporal difference error computed by the eq. 1 and the value of $st$ represents the amount of time (learning steps) the Fox agent takes to reach the best solution. Indeed, this equation is calculated on the last quarter of episodes from the entire episodes to ensure the stability of the learning process, so $i=episodes-episodes/4$. For example, if the number of episodes = 200, the fitness calculation starts from the episode 150.
The QF-tuner employed the FOX with Q-learning algorithms in a new model that automates the tuning of hyperparameters without human intervention. Setting up the QF-tuner involves determining the number of iterations $max\_iter$, the number of Fox agents $g$, and the number of runs $n$, as illustrated in fig. 1. The FOX is employed, and its initial solution serves as the initial hyperparameters ($\alpha$ and $\gamma$) of the Q-learning algorithm. Next, The Q-learning algorithm starts learning iteratively based on the number of episodes, during which reward, learning error, and convergence time values are computed to derive the fitness value using eq. 11. The Fox agent’s solution is updated if $max\_iter$ or appropriate convergence is not reached. Otherwise, the QF-tuner starts again until the number of runs ($n$) is achieved to prevent getting stuck in the local optima. Finally, the optimal solution consists of the best set of hyperparameters ($\alpha$ and $\gamma$), providing robust and practical solutions for real-world problem-solving, without the need for manual or random tuning.
Figure 1: Flowchart of the QF-tuner
4 Results and Discussion
Experiments were conducted to evaluate the performance of the QF-tuner and compare it with several established optimization algorithms, including PSO, BA, and GA. OpenAI Gym environment control tasks were utilized during the experiments, as explained in section 3.1. The parameter settings for the QF-tuner comprised 100 iterations, 30 Fox agents, 10 runs, and 200 episodes, with 200 steps for the FrozenLake control task and 500 steps for the CartPole control task. The following tables 1, 2, 2 and 3 display the performance results using hyperparameters (Alpha $\alpha$, and Gamma $\gamma$), as well as Rewards, where the maximum value for FrozenLake is one and for CartPole is 500. The processing time (Time) was also considered to measure the speed of the optimization algorithm. The reward curves in the figs. 2 and 3 have been normalized between 0 and 1 to reduce the variance and improve visualization.
In fig. 2, the QF-tuner achieved a reward of 0.95, surpassing PSO (0.78), BA (0.65), GA (0.62), and Random (0.34). table 1 shows that QF-tuner reduced learning time (steps) by 33% compared to GA, 25% compared to PSO, and 20% compared to BA. Furthermore, the average increase in the reward for QF-tuner was about 36% compared to the other optimization algorithms, while the average reduction in learning time was around 26%, demonstrating its superior performance in the FrozenLake control task.
Table 1: Empirical results using the QF-tuner and other optimization algorithms on FrozenLake control task.
Method Alpha Gamma Reward Time (s)
QF-tuner 0.74 0.97 0.95 156
PSO 1.00 0.38 0.78 210
BA 0.94 0.64 0.65 197
GA 0.34 0.83 0.62 234
Random 0.09 0.74 0.34 -
Figure 2: Rewards of the QF-tuner and other optimization algorithms on the FrozenLake control task
In fig. 3, the QF-tuner achieved a reward of 499.08, surpassing PSO (398.99), GA (215.03), Random (129.46), and BA (96.65). table 2 shows that QF-tuner reduced learning time (steps) by 26% compared to GA, 20% compared to PSO, and 14% compared to BA. Moreover, the average gain in the reward for QF-tuner was approximately 57% compared to the other optimization algorithms, while the average decrease in learning time was about 20%, showing its superior performance in the CartPole control task.
Table 2: Empirical results using the QF-tuner and other optimization algorithms on CartPole control task.
Method Alpha Gamma Reward Time (s)
QF-tuner 0.83 0.95 499.08 314
PSO 1.00 0.67 398.99 394
GA 0.84 0.71 215.03 428
Random 0.56 0.38 129.46 -
BA 0.99 0.99 96.65 367
Figure 3: Rewards of the QF-tuner and other optimization algorithms on the CartPole control task
Additionally, QF-tuner outperformed the results of [15] by 8% in reward on the $4\times 4$ FrozenLake control task. Moreover, QF-tuner achieved a 6% higher reward and a 26% reduction in learning time compared to the study in [16] on the same task. On the other hand, QF-tuner outperformed the results of [12] by 25% in reward on the CartPole control task. Thus, these results indicate that the QF-tuner improves Q-learning performance in the FrozenLake and CartPole control tasks. The improvement in reward highlights the critical role of the proposed objective function in eq. 11 in boosting the optimization process. Furthermore, the significant reduction in learning time emphasizes the effective integration of FOX with Q-learning, leading to superior performance due to their strong compatibility. However, a key limitation of the QF-tuner is its resource-intensive nature, which prevents it from being directly applied to real-world problems. Instead, it needs to be executed in a simulated environment to fine-tune the hyperparameters, which can then be utilized directly.
5 Conclusion
In conclusion, this paper presents a novel QF-tuner method for automatic hyperparameter tuning in the Q-learning algorithm using the FOX optimization algorithm (FOX). Additionally, we propose a new objective function that combines reward, learning error, and learning time into a single function, with more emphasis on the reward. The FOX utilizes the proposed objective function to minimize the fitness value derived by executing the Q-learning algorithm. The QF-tuner has been evaluated using benchmark control tasks from the OpenAI Gym, allowing for comparisons with other methods and studies. The superior results indicate that the QF-tuner can be a feasible alternative to established hyperparameter tuning methods to solve control tasks. However, tuning hyperparameters is time-intensive and requires multiple optimization iterations to reach the best set, which restricts its use in simulation environments. This limitation presents an opportunity for future research to enhance the methodology, making it applicable to real-world applications. Furthermore, other optimizers could be utilized, or further modifications could be proposed to develop alternative variations with better performance.